# On the Weierstraß form of infinite dimensional differential algebraic equations

### v1

@@ 5 Weierstraß form and solutions of port-Hamiltonian DAEs
Let $E$, $Q$ and $A$ be defined as before, such that $(E,AQ)$ defines a pH-DAE and assume, that there exists a $\omega>0$ with $\mathbb{C}_{{\rm Re}>\omega}\subseteq\rho(E,AQ)$. Then, for every $x_{0}\in{\rm ran}\,((\mu E-AQ)^{-1}E)^{p_{\mathrm{c,res}}^{(E,AQ)}+2}$ there exists a unique solution $x\colon\mathbb{R}_{t\geq 0}\to\mathcal{X}$ of the following algebraic equation
@@
The main problem with a transformation like that is, that (39) and (40) are not necessarily pH-DAEs anymore. Thus, the resulting question is under which circumstances $(AQ)_{1}E_{1}^{-1}$ and $E_{1}^{-1}(AQ)_{1}$ generate a $C_{0}$-semigroup. A possible condition for that is $AQ$ being strongly $(E,p)$-radial [18, Thm. 2.6.1].
@@
In order to show a), assume that $Q^{\ast}(\mathcal{Z}^{1})=Q^{\ast}(E(\mathcal{X}^{1}))=\mathcal{X}^{1}$ holds. Thus, $Q^{\ast}E\colon\mathcal{X}^{1}\to\mathcal{X}^{1}$ is non-negative and self-adjoint. In fact, since $Q^{\ast}E$ self-adjoint the mappings $Q^{\ast}E\pm I_{\mathcal{X}}$ are surjective. Hence, for every $z\in\mathcal{X}^{1}$ there are $x_{\pm}=x_{\pm}^{1}+x_{\pm}^{2}\in\mathcal{X}=\mathcal{X}^{1}\oplus\mathcal{X}^{2}$ with $(Q^{\ast}E\pm I_{\mathcal{X}})x_{\pm}=z$. Together with $Q^{\ast}E(\mathcal{X}^{1})=\mathcal{X}^{1}$ one has $PQ^{\ast}Ex_{\pm}=PQ^{\ast}Ex_{\pm}^{1}+PQ^{\ast}Ex_{\pm}^{2}=Q^{\ast}EPx_{\pm}^{1}=Q^{\ast}EPx_{\pm}$ and
@@
$=PQ^{\ast}Ex_{\pm}\pm Px_{\pm}$
$=Q^{\ast}EPx_{\pm}\pm Px_{\pm}=(Q^{\ast}E\pm I_{\mathcal{X}^{1}})x_{\pm}.$
Thus, $Q^{\ast}E\pm I_{\mathcal{X}^{1}}\colon\mathcal{X}^{1}\to\mathcal{X}^{1}$ are surjective and consequently, as a closed operator, $(Q^{\ast}E)|_{\mathcal{X}^{1}}$ is self-adjoint. Furthermore, since $Q^{\ast}$ and $E_{1}$ are invertible and $(Q^{\ast}E)|_{\mathcal{X}^{1}}$ is self-adjoint, $(Q^{\ast}E)|_{\mathcal{X}^{1}}$ becomes invertible and $((Q^{\ast}E)|_{\mathcal{X}^{1}})^{-1}$ self-adjoint as well. Hence, using Lemma 4.2 for $B=((Q^{\ast}E)|_{\mathcal{X}^{1}})^{-1}$ there exists a $c>0$, such that $T=((Q^{\ast}E)|_{\mathcal{X}^{1}})\geq cI_{\mathcal{X}^{1}}$ and $\langle\cdot,\cdot\rangle_{\tilde{\mathcal{X}}^{1}}\coloneqq\langle(Q^{\ast}E)|_{\mathcal{X}^{1}}\cdot,\cdot\rangle$ induces a norm equivalent to $\|\cdot\|_{\mathcal{X}}$ on $\mathcal{X}^{1}$. Define $\tilde{\mathcal{X}}^{1}\coloneqq(\mathcal{X}^{1},\langle\cdot,\cdot\rangle_{\tilde{\mathcal{X}}^{1}})$. Since $A$ is dissipative one derives
@@
which means that $E_{1}^{-1}(AQ)_{1}$ is dissipative in $\tilde{\mathcal{X}}^{1}$. Given that $\mathcal{X}=\mathcal{X}^{1}\oplus\mathcal{X}^{2}$, $E_{1}$ is bijective and $(\lambda E_{1}-(AQ)_{1})$ is surjective, $(\lambda I_{\mathcal{X}^{1}}-E_{1}^{-1}(AQ)_{1})=E_{1}^{-1}(\lambda E_{1}-(AQ)_{1})\colon\mathcal{D}((AQ)_{1})\subseteq\mathcal{X}^{1}\to\mathcal{X}^{1}$ becomes surjective as well. In the next step one shows, that this also holds for $\lambda=1$. Let $z\in{\rm ran}\,(I_{\mathcal{X}^{1}}-E_{1}^{-1}(AQ)_{1})^{\perp}$. Then, there exists a $x\in\mathcal{X}^{1}$ with $(\lambda I_{\mathcal{X}^{1}}-E_{1}^{-1}(AQ)_{1})x=z$ and
00 0 $=\langle z,(I_{\mathcal{X}^{1}}-E_{1}^{-1}(AQ)_{1})x\rangle_{\tilde{\mathcal{X}}^{1}}$
$=\langle(\lambda I_{\mathcal{X}^{1}}-E_{1}^{-1}(AQ)_{1})x,(I_{\mathcal{X}^{1}}-E_{1}^{-1}(AQ)_{1})x\rangle_{\tilde{\mathcal{X}}^{1}}$
$=\lambda\|x\|_{\tilde{\mathcal{X}}^{1}}^{2}-\langle E_{1}^{-1}(AQ)_{1}x,x\rangle_{\tilde{\mathcal{X}}^{1}}-\langle x,E_{1}^{-1}(AQ)_{1}x\rangle_{\tilde{\mathcal{X}}^{1}}+\|E_{1}^{-1}(AQ)_{1}x\|_{\tilde{\mathcal{X}}^{1}}^{2}.$
Taking the real part of this and using the dissipativity of $E_{1}^{-1}(AQ)_{1}$ on $\tilde{\mathcal{X}}^{1}$ one derives $x=0$ and $z=(\lambda I_{\mathcal{X}^{1}}-E_{1}^{-1}(AQ)_{1})x=0$. Thus, the range of $I_{\mathcal{X}^{1}}-E_{1}^{-1}(AQ)_{1}$ is dense on $\tilde{\mathcal{X}}^{1}$ and by [7, Lem. 6.1.6] ${\rm ran}\,(I_{\mathcal{X}^{1}}-E_{1}^{-1}(AQ)_{1})=\tilde{\mathcal{X}}^{1}$. By the Theorem of Lumer-Phillips [7, Thm. 6.1.7] $E_{1}^{-1}(AQ)_{1}$ generates a contraction semigroup on $\tilde{\mathcal{X}}^{1}$. In this case, $E_{1}^{-1}(AQ)_{1}$ also generates a $C_{0}$-semigroup on $\mathcal{X}^{1}$.
@@
Let $\mathcal{X}=\mathcal{Z}$ and $Q=I_{\mathcal{X}}$. Then, $E$ commutes with $A$ on $\mathcal{D}(A)$ and
@@
Since ${\rm ker}\,E\subseteq\mathcal{X}^{2}$ and $\mathcal{X}=\mathcal{X}^{1}\oplus\mathcal{X}^{2}$, one has $\mathcal{E}_{1}=I_{\mathcal{X}^{1}}$. Thus, one has $\mathcal{E}(\mathcal{X}^{1})=\mathcal{X}^{1}=\mathcal{Z}^{1}=\mathcal{E}_{1}(\mathcal{Z}^{1})$. But, in this case it is even possible to say directly something about the generation of $C_{0}$-semigroups without using Theorem 5.3. Because in such a case $\mathcal{A}_{1}=\mathcal{A}_{1}\mathcal{E}_{1}^{-1}=\mathcal{E}_{1}^{-1}\mathcal{A}_{1}$ is dissipative (as a restriction of a dissipative operator) and $\lambda I_{\mathcal{X}^{1}}-\mathcal{A}_{1}$ is surjective (since $(\omega,\infty)\subseteq\rho(\mathcal{E},\mathcal{A})=\rho(I_{\mathcal{X}^{1}},\mathcal{A}_{1})$). Hence, using the Lumer-Phillips Theorem one obtains the same outcome. This result is similar to [6, Section IV].

### v2

@@ 5 Weierstraß form and solutions of port-Hamiltonian DAEs
Let $E$, $Q$ and $A$ be defined as before, such that $(E,AQ)$ defines a pH-DAE and assume, that there exists a $\omega>0$ with $\mathbb{C}_{{\rm Re}>\omega}\subseteq\rho(E,AQ)$. Then, for every $x_{0}\in{\rm ran}\,((\mu E-AQ)^{-1}E)^{p_{\mathrm{c,res}}^{(E,AQ)}+2}$ there exists a unique solution $x\colon\mathbb{R}_{\geq 0}\to\mathcal{X}$ of the following differential algebraic equation
@@
The main problem with a transformation like that is, that (41) and (42) are not necessarily pH-DAEs anymore. Thus, the resulting question is under which circumstances $(AQ)_{1}E_{1}^{-1}$ and $E_{1}^{-1}(AQ)_{1}$ generate a $C_{0}$-semigroup. A possible condition for that is $AQ$ being strongly $(E,p)$-radial [19, Thm. 2.6.1].
@@
In order to show a), assume that $Q^{\ast}(\mathcal{Z}^{1})=Q^{\ast}(E(\mathcal{X}^{1}))=\mathcal{X}^{1}$ holds. Thus, $Q^{\ast}E\colon\mathcal{X}^{1}\to\mathcal{X}^{1}$ is non-negative and self-adjoint. In fact, since $Q^{\ast}E$ is self-adjoint the mappings $Q^{\ast}E\pm\mathrm{i}I_{\mathcal{X}}$ are surjective. Hence, for every $z\in\mathcal{X}^{1}$ there are $x_{\pm}=x_{\pm}^{1}+x_{\pm}^{2}\in\mathcal{X}=\mathcal{X}^{1}\oplus\mathcal{X}^{2}$ with $(Q^{\ast}E\pm\mathrm{i}I_{\mathcal{X}})x_{\pm}=z$. Together with $Q^{\ast}E(\mathcal{X}^{1})=\mathcal{X}^{1}$ one has $PQ^{\ast}Ex_{\pm}=PQ^{\ast}Ex_{\pm}^{1}+PQ^{\ast}Ex_{\pm}^{2}=Q^{\ast}EPx_{\pm}^{1}=Q^{\ast}EPx_{\pm}$ and
@@
$=PQ^{\ast}Ex_{\pm}\pm\mathrm{i}Px_{\pm}$
$=Q^{\ast}EPx_{\pm}\pm\mathrm{i}Px_{\pm}=(Q^{\ast}E\pm\mathrm{i}I_{\mathcal{X}^{1}})x_{\pm}.$
Thus, $Q^{\ast}E\pm\mathrm{i}I_{\mathcal{X}^{1}}\colon\mathcal{X}^{1}\to\mathcal{X}^{1}$ are surjective and consequently, as a closed operator, $(Q^{\ast}E)|_{\mathcal{X}^{1}}$ is self-adjoint. Furthermore, since $Q^{\ast}$ and $E_{1}$ are invertible and $(Q^{\ast}E)|_{\mathcal{X}^{1}}$ is self-adjoint, $(Q^{\ast}E)|_{\mathcal{X}^{1}}$ becomes invertible and $((Q^{\ast}E)|_{\mathcal{X}^{1}})^{-1}$ self-adjoint as well. Hence, using Lemma 4.2 for $B=((Q^{\ast}E)|_{\mathcal{X}^{1}})^{-1}$ there exists a $c>0$, such that $T=((Q^{\ast}E)|_{\mathcal{X}^{1}})\geq cI_{\mathcal{X}^{1}}$ and $\langle\cdot,\cdot\rangle_{\tilde{\mathcal{X}}^{1}}\coloneqq\langle(Q^{\ast}E)|_{\mathcal{X}^{1}}\cdot,\cdot\rangle$ induces a norm equivalent to $\|\cdot\|_{\mathcal{X}}$ on $\mathcal{X}^{1}$. Define $\tilde{\mathcal{X}}^{1}\coloneqq(\mathcal{X}^{1},\langle\cdot,\cdot\rangle_{\tilde{\mathcal{X}}^{1}})$. Since $A$ is dissipative one derives
@@
which means that $E_{1}^{-1}(AQ)_{1}$ is dissipative in $\tilde{\mathcal{X}}^{1}$. Given that $\mathcal{X}=\mathcal{X}^{1}\oplus\mathcal{X}^{2}$, $E_{1}$ is bijective and $(\lambda E_{1}-(AQ)_{1})$ is surjective, $(\lambda I_{\mathcal{X}^{1}}-E_{1}^{-1}(AQ)_{1})=E_{1}^{-1}(\lambda E_{1}-(AQ)_{1})\colon\mathcal{D}((AQ)_{1})\subseteq\mathcal{X}^{1}\to\mathcal{X}^{1}$ becomes surjective as well. By the Theorem of Lumer-Phillips [1, Thm.3.4.5] $E_{1}^{-1}(AQ)_{1}$ generates a contraction semigroup on $\tilde{\mathcal{X}}^{1}$. In this case, $E_{1}^{-1}(AQ)_{1}$ also generates a $C_{0}$-semigroup on $\mathcal{X}^{1}$.
@@
Let $\mathcal{X}=\mathcal{Z}$ and $Q=I_{\mathcal{X}}$ and assume that $E$ commutes with $A$ on $\mathcal{D}(A)$. Thus
@@
Since ${\rm ker}\,E\subseteq\mathcal{X}^{2}$ and $\mathcal{X}=\mathcal{X}^{1}\oplus\mathcal{X}^{2}$, one has $\mathcal{E}_{1}=I_{\mathcal{X}^{1}}$. Thus, one has $\mathcal{E}(\mathcal{X}^{1})=\mathcal{X}^{1}=\mathcal{Z}^{1}=\mathcal{E}_{1}(\mathcal{Z}^{1})$. But, in this case it is even possible to say directly something about the generation of $C_{0}$-semigroups without using Theorem 5.3. Because in such a case $\mathcal{A}_{1}=\mathcal{A}_{1}\mathcal{E}_{1}^{-1}=\mathcal{E}_{1}^{-1}\mathcal{A}_{1}$ is dissipative (as a restriction of a dissipative operator) and $\lambda I_{\mathcal{X}^{1}}-\mathcal{A}_{1}$ is surjective (since $(\omega,\infty)\subseteq\rho(\mathcal{E},\mathcal{A})=\rho(I_{\mathcal{X}^{1}},\mathcal{A}_{1})$). Hence, using the Lumer-Phillips Theorem one obtains the same outcome. This result is similar to [7, Section IV].
@@ Acknowledgements
Data Availability No data was used for the research described in the article.
Declarations
Conflicts of interest On behalf of all authors, the corresponding author states that there is no conflict of interest.